# score-based generative model SNR signal-to-noise ratio

### v2

@@ IV-C DNN configuration and training
For Stream.FM, we parameterize our architecture described in Section III-C with two residual blocks per level and four U-Net levels with [128, 256, 256, 256] channels, respectively, leading to 27.9 M parameters when used as an FM backbone DNN. For the initial predictor network $D_{\phi}$ in the SE task, we remove all time-conditioning layers and reduce the complex input channels from two to one, resulting in 24.6 M parameters. For the non-causal FM baseline models, we use the original NCSN++ architecture [song2021scorebased], here also parameterized with four U-Net levels in the same channel configuration and also two residual blocks per level (38.7 M parameters).
For each task, we train an FM model $v_{\theta}$ using Eq. 4 for 150,000 steps, 2 GPUs, and 2-second random snippets with a batch size of 12 per GPU using the SOAP optimizer [vyas2025soap] which was recently found to perform well for diffusion model training [schaipp2025optimization].We use a cosine annealing learning rate schedule with a maximum learning rate of $\lambda=5\times 10^{-4}$ and linear warmup for the first 1,000 steps, clamping the scheduled $\lambda$ to a minimum value of $10^{-6}$. We use gradient clipping with a maximum norm of $\lVert\nabla\rVert_{\max}=3$ for all tasks except for SE with $\lVert\nabla\rVert_{\max}=1$ and codec post-filtering with $\lVert\nabla\rVert_{\max}=5$, based on empirical gradient norm inspection.
To train the SE model, we first train the initial predictor $D_{\phi}$ for 150,000 steps using Eq. 17 with the SOAP optimizer [vyas2025soap] at a constant learning rate of $\lambda=3\times 10^{-3}$. We then freeze the initial predictor during FM model training. For SE, we also train a lower-latency joint predictive-generative Stream.FM model, reconfiguring the STFT with 16 ms frames and 8 ms hops for a total latency of only 24 ms, keeping all other settings the same.
@@ IV-D1 Baseline methods
As streaming-capable baseline methods for SE, we use DEMUCS [defossez2020realtime], DeepFilterNet3 [schroeter2023deepfilternet3], HiFi-Stream [dmitrieva2025hifi], and Diffusion Buffer [lay2025diffusionjpre]. We further include the aTENNuate [pei2025optimized] method for which no streaming model variant has been published, hence we only evaluate its offline performance. We use the official streaming implementations for DEMUCS [^5] https://github.com/facebookresearch/denoiser, master64 checkpoint., HiFi-Stream [^6] https://github.com/KVDmitrieva/source_sep_hifi, hifi_fms checkpoint. and DeepFilterNet3777https://github.com/Rikorose/DeepFilterNet, and the official Python package for aTENNuate [^8] https://pypi.org/project/attenuate/.
@@ IV-D3 Metric evaluation
As intrusive metrics, we report wideband PESQ [Rix2001PESQ], ESTOI [jensen2016estoi], SI-SDR [roux2019sdr] and LSD999as in https://github.com/haoheliu/ssr_eval, 32 ms Hann window, 75% overlap. [liu2022neural]. We further report the word error rate (WER) using the QuartzNet15x5Base-En model [kriman2020quartznet] from the NeMo toolkit [kuchaiev2019nemo] as the speech recognition backend, using the model’s transcripts of the clean audios as the reference. As non-intrusive metrics, we report NISQA [mittag2021nisqa], WVMOS [andreev2023hifipp], and DistillMOS [stahl2025distillmos] which we refer to as DiMOS for brevity.
@@ IV-D4 Listening experiments
We conduct two MUSHRA-like listening experiments [mushra2014method], one for SE and one for BWE, each with 12 participants who gave informed consent. We asked participants to rate the overall quality (0–100) of 8 randomly sampled utterances, as reconstructed by each method. For SE, we compare predictive-generative Stream.FM (SFM) against non-streaming FM, both with 4 Euler steps, Diffusion Buffer [lay2025diffusionjpre] at $d\in\{0,9\}$, and DEMUCS [defossez2020realtime]. For BWE, we compare SFM against non-streaming FM, both with 5 Euler steps, against SFM with a learned RK5 solver. We use the noisy/downsampled utterance as the low anchor for SE/BWE, respectively.
@@ V Results and discussion
Figure 2: Violin plots of the scores listeners assigned to examples from each method in the listening experiments for (a) speech enhancement and (b) bandwidth extension. SFM is Stream.FM, FM is the flow matching baseline, and DB is the Diffusion Buffer [lay2025diffusionjpre]. Note that FM has $\ell_{\text{alg}}=\infty$ and DB with $d=9$ has $\ell_{\text{alg}}\approx 180$ ms.
TABLE II: Mean metrics for speech enhancement on EARS-WHAM v2 (16 kHz) [richter2024ears]. Our main model (SFM) is evaluated using different ODE solvers (Euler1, Euler4, etc.) with the number indicating the number of solver steps, and compared against several streaming and non-streaming baselines. “LRK4” refers to a four-stage learned Runge-Kutta solver, see Section III-E. In the NFE column, 1+n indicates a single call for the initial predictor and $n$ calls for the flow model. Methods marked with use author-provided model checkpoints trained on different data. $\ell_{\text{alg}}$ is the algorithmic latency, see Section IV-D2. Best within a group in bold, second best underlined, worse than input in red.
@@
SFM Euler1 1+1 2.18 0.84 15.2 3.91 2.65 4.43 19.5% 1.39 32 52.5M
SFM Euler4 1+4 2.09 0.83 14.3 3.88 2.72 4.50 21.8% 1.29 32 52.5M
SFM Midpoint2 1+4 2.02 0.82 13.3 3.66 2.72 4.28 23.6% 1.38 32 52.5M
SFM LRK4 1+4 2.24 0.82 14.0 3.81 3.00 4.06 20.2% 1.42 32 52.5M
@@
Diffusion Buffer $d=0$ [lay2025diffusionjpre] 1 1.75 0.75 11.1 2.81 2.08 2.49 27.3% 1.37 32 22.2M
Diffusion Buffer $d=9$ [lay2025diffusionjpre] 1 2.06 0.81 14.4 3.64 2.44 3.78 21.7% 1.47 176 22.2M
DEMUCS [defossez2020realtime] 1 1.95 0.79 13.0 3.46 3.00 2.61 27.1% 1.55 41 33.5M
DeepFilterNet3 [schroeter2023deepfilternet3] 1 1.76 0.70 8.8 3.29 2.81 3.37 37.0% 2.46 40 2.14M
HiFi-Stream [dmitrieva2025hifi] 1 1.21 0.37 -3.8 1.94 1.49 1.50 69.7% 1.57 256 1.6M
@@
FM Euler1 1+1 2.36 0.86 16.8 4.20 2.73 4.37 16.4% 1.53 $\infty$ 73.7M
FM Euler4 1+4 2.41 0.86 16.1 4.34 2.82 4.50 18.4% 1.31 $\infty$ 73.7M
FM Midpoint2 1+4 2.35 0.85 15.0 4.29 2.82 4.39 19.4% 1.44 $\infty$ 73.7M
aTENNuate [pei2025optimized] 1 1.86 0.72 9.0 2.95 2.86 2.37 33.1% 1.42 $\infty$ 0.8M
TABLE III: VoiceBank-DEMAND speech enhancement benchmark [valentini2016investigating]. “DB” refers to the Diffusion Buffer [lay2025diffusionjpre].
Method PESQ ESTOI SI-SDR DiMOS $\ell_{\text{alg}}$ (ms)
@@
SFM LRK4 2.69 0.84 13.3 3.81 32
DB $d=0$ [lay2025diffusionjpre] 2.42 0.81 13.1 2.75 32
DB $d=9$ [lay2025diffusionjpre] 2.45 0.84 14.5 3.66 176
DEMUCS [defossez2020realtime] 2.60 0.85 15.1 3.46 41
DeepFilterNet3 [schroeter2023deepfilternet3] 2.71 0.84 17.3 3.34 40
HiFi-Stream [dmitrieva2025hifi] 2.48 0.83 14.1 3.35 256
aTENNuate [pei2025optimized] 2.97 0.84 16.2 2.95 $\infty$
@@ V-A Speech enhancement
We show the metric results for the SE task in Table II. We find that all SFM variants attain the highest values among the streaming-capable baselines in almost all metrics and exhibit particularly strong improvements in non-intrusive DistillMOS, WVMOS and NISQA, though WVMOS rates DEMUCS [defossez2020realtime] on-par. When the Diffusion Buffer (DB) baseline [lay2025diffusionjpre] is configured to use the same low algorithmic latency ($d=0$) as SFM, where DB uses only a single diffusion step per frame, SFM with four Euler steps shows strong advantages over DB. SFM also performs similar to or better than the higher-latency DB variant $(d=9)$ with 10 diffusion steps, the main model proposed in [lay2025diffusionjpre]. HiFi-Stream [dmitrieva2025hifi] and aTTENuate [pei2025optimized] do not achieve clear improvements over the noisy mixtures and in particular worsen the WER. This may indicate an overfitting to their respective training datasets, which can be seen as reasonable under their small DNN parameter budgets.
Comparing different ODE solvers for Stream.FM, we see that a single Euler step (Euler1) shows excellent performance in this task, but 4 Euler steps (Euler4) slightly improve most non-intrusive metrics and LSD, at some decrease in intrusive metric scores. The learned RK solver (LRK4) strongly improves PESQ, WVMOS and WER over Euler4, but does not yield a metric improvement across the board. The non-streaming FM baseline outperforms SFM, but the quality degradation of the streaming models is expected due to the small look-ahead, and is on acceptable levels. Our $\ell_{\text{alg}}=16$ ms lower-latency SFM variant exhibits a relatively minor quality reduction compared to the main SFM model, proving the viability of our method for SE in even lower-latency settings ($\ell_{\text{tot}}\approx 24$ ms). In Table III, we also show metrics on the classic VoiceBank-DEMAND benchmark [valentini2016investigating]. SFM exhibits the generalization capabilities expected of diffusion-based SE models [richter2023speech], and attains the best PESQ, ESTOI and DistillMOS among streaming methods.
@@ V-B Dereverberation
TABLE IV: Dereverberation task (EARS-Reverb v2 test set). SFM has $\ell_{\text{alg}}=32$ ms, $\ell_{\text{tot}}\approx 48$ ms; FM has $\ell_{\text{alg}}=\infty$ and 38.7M parameters. Best within a group in bold, second best underlined, worse than input in red.
@@
Reverberant - 1.32 0.58 -16.6 3.02 2.02 2.11 20.1% 1.21
SFM Euler1 1 1.63 0.73 -14.2 3.29 2.17 3.03 23.6% 1.49
SFM Euler5 5 2.01 0.79 -13.3 3.76 2.49 3.59 16.8% 1.12
SFM Midpoint2 4 1.94 0.78 -14.3 3.70 2.35 3.62 17.3% 0.98
SFM LRK5 5 1.91 0.78 -13.5 3.49 2.44 3.29 16.6% 1.01
FM Euler5 5 2.31 0.85 -11.7 3.77 2.43 3.47 11.4% 1.01
TABLE V: Lyra V2 codec post-filtering task on EARS-WHAM v2 test set clean utterances. SFM has $\ell_{\text{alg}}=40$ ms, $\ell_{\text{tot}}\approx 60$ ms and 27.9M parameters; FM has $\ell_{\text{alg}}=\infty$ and 38.7M parameters. Best within a group in bold, second best underlined, worse than input in red.
@@
Decoded - 2.00 0.76 1.6 3.08 2.60 2.68 13.6% 1.01
SFM Euler1 1 1.80 0.63 4.9 2.18 1.81 2.45 44.3% 5.84
SFM Euler5 5 2.55 0.80 3.1 4.00 2.93 3.96 16.3% 1.49
SFM Midpoint2 4 2.38 0.79 1.9 4.09 2.80 4.11 16.0% 1.25
SFM LRK5 5 2.27 0.77 1.8 3.94 2.71 3.90 16.3% 1.09
FM Euler5 5 2.56 0.80 3.0 4.14 2.92 3.97 16.1% 1.41
TABLE VI: Bandwidth extension task ($\{2,4\}$ kHz $\to 8$ kHz of frequency content) on EARS-WHAM v2 test set clean utterances. SFM has $\ell_{\text{alg}}=32$ ms, $\ell_{\text{tot}}\approx 48$ ms and 27.9M parameters; FM has $\ell_{\text{alg}}=\infty$ and 38.7M parameters. Best within a group in bold, second best underlined, worse than input in red.
@@
Bandlimited - 3.51 0.84 15.9 3.09 2.21 2.93 19.4% 2.24
SFM Euler1 1 3.22 0.92 16.8 3.43 2.41 3.32 15.3% 1.95
SFM Euler5 5 3.37 0.94 16.5 4.07 2.96 3.76 12.3% 1.26
SFM Midpoint2 4 3.10 0.94 16.0 4.18 3.01 3.97 12.0% 1.10
SFM LRK5 5 3.02 0.94 15.3 4.19 3.02 3.99 10.5% 1.00
FM Euler5 5 3.52 0.94 16.3 4.28 3.00 3.77 11.8% 1.29
TABLE VII: STFT phase retrieval task with 50% overlap, no lookahead, on EARS-WHAM v2 test set clean utterances. RTISI-DM refers to [peer2024flexible] using the Difference Map algorithm [elser2003phaseretrieval] with hyperparameter $\beta=1.75$, which we grid-searched as the optimum for 50 iterations. RTISI has no parameters. For RTISI and RTISI-DM, NFE refers to the number of algorithm iterations. Best within a group in bold, second best underlined, worse than input in red.

### v3

@@ IV-C DNN configuration and training
For Stream.FM, we parameterize our architecture described in Section III-C with two residual blocks per level and four U-Net levels with [128, 256, 256, 256] channels, respectively, leading to 27.9 M parameters when used as an FM backbone DNN. For the initial predictor network $D_{\phi}$ in the SE task, we remove all time-conditioning layers and reduce the complex input channels from two to one, resulting in 24.6 M parameters. For the non-causal FM baseline models, we use the original NCSN++ architecture [54], here also parameterized with four U-Net levels in the same channel configuration and also two residual blocks per level (38.7 M parameters).
For each task, we train a flow model $v_{\theta}$ using Eq. 4 for 150,000 steps on two NVIDIA RTX A6000 GPUs, using 2-second random snippets with a batch size of 12 per GPU. We use the SOAP optimizer [57] which was recently found to perform well for diffusion model training [52].We use a cosine annealing learning rate schedule with a maximum learning rate of $\lambda=5\times 10^{-4}$ and linear warmup for the first 1,000 steps, clamping the scheduled $\lambda$ to a minimum value of $10^{-6}$. We use gradient clipping with a maximum norm of $\lVert\nabla\rVert_{\mathrm{max}}=3$ for all tasks except for SE with $\lVert\nabla\rVert_{\mathrm{max}}=1$ and codec post-filtering with $\lVert\nabla\rVert_{\mathrm{max}}=5$, based on empirical gradient norm inspection.
To train the SE model, we first train the initial predictor $D_{\phi}$ for 150,000 steps using Eq. 16 with the SOAP optimizer [57] at a constant learning rate of $\lambda=3\times 10^{-3}$. We then freeze the initial predictor during FM model training. For SE, we also train a lower-latency joint predictive-generative Stream.FM model, reconfiguring the STFT with 16 ms frames and 8 ms hops for a total latency of only 24 ms, keeping all other settings the same.
@@ IV-D1 Baseline methods
As streaming-capable baseline methods for SE, we use DEMUCS [9], DeepFilterNet3 [53], HiFi-Stream [10], CleanUMamba [12], Diffusion Buffer [28], and a SEMamba [7] variant modified for streaming and trained [28] on EWv2 [49]. We further include the aTENNuate [42] method for which no streaming model variant has been published, hence we only evaluate its offline variant. As offline-only baselines, we use the non-causal FM model described in Section IV-C, SBVE [21, 46], MambAttention [25], and AnyEnhance [68, 69]. We use the official streaming implementations for DEMUCS [^7] https://github.com/facebookresearch/denoiser, master64 checkpoint., HiFi-Stream [^8] https://github.com/KVDmitrieva/source_sep_hifi, hifi_fms checkpoint., DeepFilterNet3999https://github.com/Rikorose/DeepFilterNet, DeepFilterNet3 checkpoint., and CleanUMamba [^10] https://github.com/lab-emi/CleanUMamba, E8_pruned-5M checkpoint., and the official Python package for aTENNuate [^11] https://pypi.org/project/attenuate/. For SBVE, we use the checkpoint [^12] https://github.com/sp-uhh/sgmse trained on 16 kHz EARS-WHAM [49] and VB-DMD [56].
As baselines for other tasks, we use: AnyEnhance [68, 69] for dereverberation, BWE and codec artifact removal; for dereverberation, a 48 kHz SGMSE+ model [49] trained on EARS-Reverb v1; for STFT PR, DiffPhase [40]; for Mel vocoding, HiFi-GAN [22] with the 16 kHz SpeechBrain checkpoint [45].
@@ IV-D3 Metric evaluation
As intrusive metrics, we report wideband PESQ [50], ESTOI [20], SI-SDR [29] and log-spectral distance (LSD)131313as in https://github.com/haoheliu/ssr_eval, 32 ms Hann window, 75% overlap. [35]. We further report the word error rate (WER) using the QuartzNet15x5Base-En model [23] from the NeMo toolkit [24] as the speech recognition backend, using the model’s transcripts of the clean audios as the reference. As non-intrusive metrics, we report NISQA [39], WVMOS [2], and DistillMOS [55] which we refer to as DiMOS for brevity. We evaluate all model outputs at 16 kHz, downsampling to 16 kHz as necessary, e.g., for AnyEnhance [68].
@@ IV-D4 Listening experiments
We conduct two MUSHRA-like listening experiments [19], one for SE and one for BWE, each with 12 participants who gave informed consent. We asked participants to rate the overall quality (0–100) of 8 randomly sampled utterances, as reconstructed by each method. For SE, we compare predictive-generative Stream.FM (SFM) against non-streaming FM, both with 4 Euler steps, Diffusion Buffer [28] at $d\in\{0,9\}$, and DEMUCS [9]. For BWE, we compare SFM against non-streaming FM, both with 5 Euler steps, against SFM with a learned RK5 solver. We use the noisy/downsampled utterance as the low anchor for SE/BWE, respectively.
@@ V Results and discussion
Figure 2: Violin plots of the scores listeners assigned to examples from each method in the listening experiments for (a) speech enhancement and (b) bandwidth extension. SFM is Stream.FM, FM is the flow matching baseline, and DB is the Diffusion Buffer [28]. Note that FM has $\ell_{\text{alg}}=\infty$ and DB with $d=9$ has $\ell_{\text{alg}}\approx 180$ ms.
TABLE II: Mean metrics for speech enhancement on EWv2 (16 kHz) [49]. Our main model (SFM) is evaluated using different ODE solvers (Euler1, Euler4, etc.) with the number indicating the number of solver steps, and compared against several streaming and non-streaming baselines. “LRK4” refers to a four-stage learned Runge-Kutta solver, see Section III-E. In the NFE column, 1+n indicates a single call for the initial predictor and $n$ calls for the flow model. Methods marked with use author-provided model checkpoints trained on different data. Best within a group bold, second best underlined, worse than input in red. $\ell_{\text{alg}}$ and $\ell_{\text{tot}}$ in milliseconds.
@@
SFM Euler1 1+1 2.18 0.84 15.2 3.91 2.65 4.43 19.5% 1.39 32 48 52.5M
SFM Euler4 1+4 2.09 0.83 14.3 3.88 2.72 4.50 21.8% 1.29 32 48 52.5M
SFM Midpoint2 1+4 2.02 0.82 13.3 3.66 2.72 4.28 23.6% 1.38 32 48 52.5M
SFM LRK4 (18) 1+4 2.24 0.82 14.0 3.81 3.00 4.06 20.2% 1.42 32 48 52.5M
SFM LRK4 (19) 1+4 2.30 0.83 14.1 3.70 2.79 4.04 20.1% 1.36 32 48 52.5M
Initial Predictor $D_{\phi}$ 1 2.11 0.79 13.5 3.64 2.60 3.56 25.4% 1.42 32 48 24.6M
@@
Diffusion Buffer $d=0$ [28] 1 1.75 0.74 10.9 2.75 2.17 2.45 27.3% 1.45 32 48 22.2M
Diffusion Buffer $d=9$ [28] 1 2.09 0.81 15.0 3.66 2.56 3.81 21.7% 1.57 176 192 22.2M
SEMamba (causal) [7, 28] 1 2.61 0.82 8.8 3.75 2.65 3.25 18.6% 1.16 25 31 1.24M
DeepFilterNet3 [53] 1 1.76 0.70 8.8 3.29 2.81 3.37 37.0% 2.46 40 50 2.14M
DEMUCS [9] 1 1.95 0.79 13.0 3.46 3.00 2.61 27.1% 1.55 41 57 33.5M
CleanUMamba [12] 1 1.98 0.79 13.5 3.49 2.84 2.72 27.5% 1.08 48 64 4.9M
HiFi-Stream [10] 1 1.21 0.37 -3.8 1.94 1.49 1.50 69.7% 1.57 256 384 1.6M
@@
FM Euler1 1+1 2.36 0.86 16.8 4.20 2.73 4.37 16.4% 1.53 $\infty$ $\infty$ 73.7M
FM Euler4 1+4 2.41 0.86 16.1 4.34 2.82 4.50 18.4% 1.31 $\infty$ $\infty$ 73.7M
SBVE [21, 46] 60 2.15 0.84 12.9 4.39 3.14 4.18 18.4% 1.60 $\infty$ $\infty$ 65.6M
AnyEnhance (44.1 kHz) [68] 1 1.80 0.72 6.0 4.02 2.63 3.81 30.7% 1.41 $\infty$ $\infty$ 45.7M
MambAttention [25] 1 2.06 0.73 7.4 3.62 2.71 3.24 31.3% 1.17 $\infty$ $\infty$ 2.33M
aTENNuate [42] 1 1.86 0.72 9.0 2.95 2.86 2.37 33.1% 1.42 $\infty$ $\infty$ 0.8M
TABLE III: VoiceBank-DEMAND speech enhancement benchmark (16 kHz) [56]. “DB” refers to the Diffusion Buffer [28]. DB, SFM, FM, and Causal SEMamba were trained on EWv2. All models configured as in Table II. $\ell_{\text{alg}}$ in milliseconds.
Method PESQ ESTOI SI-SDR DiMOS $\ell_{\text{alg}}$
@@
SFM LRK4 (18) 2.69 0.84 13.3 3.81 32
SFM LRK4 (19) 2.72 0.84 13.0 3.70 32
DB $d=0$ [28] 2.42 0.81 13.1 2.75 32
DB $d=9$ [28] 2.45 0.84 14.5 3.66 176
SEMamba [7, 28] 3.18 0.79 5.8 3.52 31
DeepFilterNet3 [53] 2.71 0.84 17.3 3.34 40
DEMUCS [9] 2.60 0.85 15.1 3.46 41
CleanUMamba [12] 2.64 0.85 16.4 3.47 48
HiFi-Stream [10] 2.48 0.83 14.1 3.35 256
@@
SBVE [21, 46] 3.06 0.89 19.2 3.90 $\infty$
AnyEnhance [69, 68] 2.72 0.81 10.8 3.84 $\infty$
aTENNuate [42] 2.97 0.84 16.2 2.95 $\infty$
@@ V-A Speech enhancement
We show the metric results for the SE task in Table II. We find that all SFM variants attain the best or second-best values among streaming-capable methods in almost all metrics and exhibit particularly strong improvements in non-intrusive DistillMOS, WVMOS and NISQA. We also list the metrics for the outputs of the initial predictor $D_{\phi}$, which demonstrate that the subsequent flow matching stage with the LRK4 (19) solver clearly improves all metrics. When the Diffusion Buffer (DB) baseline [28] is configured to use the same low algorithmic latency ($d=0$) as SFM, where DB uses only a single diffusion step per frame, SFM with four Euler steps shows strong advantages over DB. SFM also performs similar to or better than the higher-latency DB variant $(d=9)$ with 10 diffusion steps, the main model proposed in [28].
Causal SEMamba [7, 28] achieves the best PESQ and WER values but falls short on SI-SDR and non-intrusive metrics, suggesting excellent denoising ability but suboptimal speech quality. CleanUMamba [12], while not trained on EWv2, shows the best overall LSD value and decent performance in most other metrics except NISQA. HiFi-Stream [10] and aTTENuate [42] do not achieve clear improvements over the noisy mixtures and in particular worsen the WER. This may indicate an overfitting to their respective training datasets, which can be seen as reasonable under their small DNN parameter budgets. The non-causal generative method AnyEnhance [1] has good non-intrusive metric values but falls short on all intrusive metrics, which may be related to performing enhancement in a compressed token space. SBVE [21, 46], also a non-causal generative method, shows good quality overall but is mostly outperformed by our non-causal FM model, and uses an expensive 60 NFE. In Fig. S7 of the supplementary material, we further provide a detailed metric evaluation figure split per input SNR, which reveals that all methods exhibit relatively stable behavior over the full input SNR range.
Comparing different ODE solvers for Stream.FM, we see that a single Euler step (Euler1) shows excellent performance in this task, but 4 Euler steps (Euler4) slightly improve most non-intrusive metrics and LSD, at some decrease in intrusive metric scores. The LRK solver (LRK4) with loss (18) strongly improves PESQ, WVMOS and WER over Euler4, but does not yield a metric improvement across the board. Using the PESQ-based loss (19) instead to train the solver improves PESQ, WER, and LSD, though at the cost of some decrease in non-intrusive metrics. The non-streaming FM baseline outperforms SFM, but the quality degradation of the streaming models is expected due to the small look-ahead, and is on acceptable levels. Our $\ell_{\text{alg}}=16$ ms lower-latency SFM variant shows relatively minor quality reduction compared to the main SFM model, proving the viability of our method for SE in even lower-latency settings ($\ell_{\text{tot}}\approx 24$ ms).
In Table III, we show metrics on the classic VB-DMD benchmark [56]. SFM exhibits the generalization capabilities expected of diffusion-based SE models [47] and attains the best ESTOI and DistillMOS among streaming methods. SFM is second-best in PESQ, outperformed only by causal SEMamba [7] which however degrades SI-SDR below the noisy input and does not improve ESTOI. Both LRK solver variants, which were trained on EWv2 data, exhibit reasonable cross-dataset transfer and perform similarly to Euler4.
@@ V-B Dereverberation
TABLE IV: Metric evaluations for the restoration tasks of dereverberation, codec post-filtering, and bandwidth extension. Best within a group bold, second best underlined, worse than input in red. $\ell_{\text{alg}}$ and $\ell_{\text{tot}}$ in milliseconds.
(a) Dereverberation task (EARS-Reverb v2 test set). SGMSE+ [47] was evaluated with the official checkpoint trained on EARS-Reverb_v1 [49] in 48 kHz, and AnyEnhance [68] uses the official checkpoint with inputs resampled to 44.1 kHz. Both baselines were evaluated using outputs resampled to 16 kHz.
@@
Reverberant - 1.32 0.58 -16.6 3.02 2.02 2.11 20.1% 1.21 - - -
SFM Euler1 1 1.63 0.73 -14.2 3.29 2.17 3.03 23.6% 1.49 32 48 27.9M
SFM Euler5 5 2.01 0.79 -13.3 3.76 2.49 3.59 16.8% 1.12 32 48 27.9M
SFM Midpoint2 4 1.94 0.78 -14.3 3.70 2.35 3.62 17.3% 0.98 32 48 27.9M
SFM LRK5 (18) 5 1.91 0.78 -13.5 3.49 2.44 3.29 16.6% 1.01 32 48 27.9M
SFM LRK5 (19) 5 2.05 0.79 -13.5 3.68 2.48 3.67 15.9% 1.05 32 48 27.9M
FM Euler5 5 2.31 0.85 -11.7 3.77 2.43 3.47 11.4% 1.01 $\infty$ $\infty$ 38.7M
SGMSE+ (48 kHz) [49] 60 1.95 0.77 -15.5 3.47 2.37 3.53 15.2% 1.11 $\infty$ $\infty$ 64.7M
AnyEnhance (44.1 kHz) [69, 68] 1 1.53 0.65 -17.1 3.15 2.10 2.84 27.1% 1.46 $\infty$ $\infty$ 45.7M
(b) Lyra V2 codec post-filtering task on EWv2 test set clean utterances.
@@
Decoded - 2.00 0.76 1.6 3.08 2.60 2.68 13.6% 1.01 - - -
SFM Euler1 1 1.80 0.63 4.9 2.18 1.81 2.45 44.3% 5.84 40 60 27.9M
SFM Euler5 5 2.55 0.80 3.1 4.00 2.93 3.96 16.3% 1.49 40 60 27.9M
SFM Midpoint2 4 2.38 0.79 1.9 4.09 2.80 4.11 16.0% 1.25 40 60 27.9M
SFM LRK5 (18) 5 2.27 0.77 1.8 3.94 2.71 3.90 16.3% 1.09 40 60 27.9M
SFM LRK5 (19) 5 2.58 0.80 3.9 3.83 2.97 3.74 17.0% 1.68 40 60 27.9M
FM Euler5 5 2.56 0.80 3.0 4.14 2.92 3.97 16.1% 1.41 $\infty$ $\infty$ 38.7M
AnyEnhance (16 kHz) [68, 69] 1 1.90 0.74 0.5 3.76 2.62 3.53 18.4% 1.26 $\infty$ $\infty$ 45.7M
(c) Bandwidth extension task ($\{2,4\}$ kHz $\to 8$ kHz of frequency content) on EWv2 test set clean utterances.
@@
Bandlimited - 3.51 0.84 15.9 3.09 2.21 2.93 19.4% 2.24 - - -
SFM Euler1 1 3.22 0.92 16.8 3.43 2.41 3.32 15.3% 1.95 32 48 27.9M
SFM Euler5 5 3.37 0.94 16.5 4.07 2.96 3.76 12.3% 1.26 32 48 27.9M
SFM Midpoint2 4 3.10 0.94 16.0 4.18 3.01 3.97 12.0% 1.10 32 48 27.9M
SFM LRK5 (18) 5 3.02 0.94 15.3 4.19 3.02 3.99 10.5% 1.00 32 48 27.9M
SFM LRK5 (19) 5 3.71 0.94 16.8 3.89 2.81 3.46 11.5% 1.45 32 48 27.9M
FM Euler5 5 3.52 0.94 16.3 4.28 3.00 3.77 11.8% 1.29 $\infty$ $\infty$ 38.7M
AnyEnhance (44.1 kHz) [69, 68] 1 2.49 0.85 6.9 4.33 3.19 4.13 18.1% 1.37 $\infty$ $\infty$ 45.7M